-3/5[2/3-1/2(3/4-1/2)]
I have worked this equation to death and come up with -3/5 but the book says the answer is -13/40.....please help me see how they arrived at this answer. Thanks!!!

Answer by sdmmadam@yahoo.com(530)

-3/5[2/3-1/2(3/4-1/2)]
=-3/5[2/3-(1/2)X(3/4)+(1/2)X(1/2)]
=-3/5[2/3-3/8+1/4]
=-3/5[(2X8)-(3X3)+(1X6)]/24 (the lcm of 3,8 and 4 is 24 and hence multiplying every nr and dr of every term by 24)
=(-3/5)X[(16-9+6)/24]
=(-3/5)X[13/(3X8)]
=(-13)/40 (canceling 3)(and multiplying 5 and 8 in the dr.
Answer:(-13/40)